# for the Hall C SIDIS Collaboration

### v1

@@ IV.4 Pion multiplicities near $\phi^{*}=180^{\circ}$
Figure 19: Same as Fig. 17 but now for the experimental multiplicities averaged over $150<\phi^{*}<210$ degrees.
Due to the experimental setup limitations, it was not possible to obtain full azimuthal coverage at large $P_{t}$. Nonetheless, a considerable amount of time was spent accumulating data near $\phi^{*}=$180, which primarily measures $M_{0}(1-A+B)$. Given this caveat, it is of interest to plot the multiplicities at $\left<\phi^{*}\right$$\sim$ 180, over the full $P_{t}$ experimental range, as shown in Fig. 19. The MAPS calculations for $M_{0}$ (MAPS did not include non-zero values of A and B in their fits) describe the $P_{t}$-dependence quite well up to $P_{t}=0.5$ GeV, which is the largest value provided by the collaboration. Although not perfect, the Gaussian-shaped fits shown in the figure describe the $P_{t}$-dependence remarkably well to values of $P_{t}$ as high as 0.7 GeV.
@@
Figure 21: Ratios to $(p,\pi^{+})$ of experimental multiplicities averaged over $150<\phi^{*}<210$ degrees for $(d,\pi^{+})$ (red), $(p,\pi^{-})$ (green), and $(d,\pi^{-})$ (blue). The solid curves are the ratios of MAPS calculations, while the dashed curves are the ratios of Gaussian fits from Fig. 19.
To further quantify the observed flavor independence of the multiplicities, the ratio of $\left<\phi^{*}\right$$\sim$ 180 multiplicities for $(d,\pi^{+})$, $(p,\pi^{-})$, and $(d,\pi^{-})$ to the multiplicities for $(p,\pi^{+})$ are plotted in Fig. 21, as a function of $P_{t}$ for the three kinematic settings and four $z$ bins. The ratios are consistent with no statistically significant dependence on $P_{t}$. The ratios are compared to those from the MAPS calculations Bacchetta et al. (2022, 2024); Cerutti , which also show only very slight $P_{t}$-dependence. As expected from the discussion on normalization factors ($k$ parameter) in Sec. IV.3, the ratios for $(p,\pi^{+})$ are larger for the data than for the calculations, especially for the lower-$W$ kinematic settings. Also plotted are the ratios of the Gaussian fits displayed in Fig. 19.
@@ IV.5 Azimuthal dependence of multiplicities
The $\phi^{*}$ dependence of the measured multiplicities is quantified by the two coefficients, $A$ and $B$, associated with the $\cos(\phi^{*})$ and $\cos{(2\phi^{*})}$ modulations of the multiplicities. The $\cos(\phi^{*})$ coefficient, $A$, obtained from the fit of the multiplicity results in each $(z,P_{t})$ bin to functional form– Eq. 10, is shown in Fig. 22 as a function of $P_{t}$. These results show an overall trend that $A$ for $\pi^{-}$ production on both protons and deuterons is significantly >> 0 at high $z$ for all three kinematic settings. On the other hand, the $A$ coefficient is consistent with zero within experimental uncertainties or has small negative values for $\pi^{+}$ production on both protons and deuterons. The $\pi^{+}$ results are consistent with the previous HERMES measurements  Airapetian et al. (2013b) but have the opposite sign for the $\pi^{-}$, except for the highest $x$ HERMES point. We note that all of these data generally lie at higher $x$ than those from HERMES or COMPASS.
Figure 22: Results for the coefficients AA that scale the $\cos(\phi^{*})$ distributions, in the same format as Fig. 17. The curves are from a fit Barone et al. (2015) to HERMES data Airapetian et al. (2013b).
Similarly the $\cos{(2\phi^{*})}$ coefficient, $B$, obtained from the fit of the multiplicity results in each $(z,P_{t})$ bin to Eq. 10 is shown in Fig. 23 as a function of $P_{t}$. Other than a couple of $z$ bins, these results show small values of $B$, that are either consistent with zero or have small positive values. These results are consistent with the previous HERMES measurements Airapetian et al. (2013b) except at $\left<z\right$>=0.55.
Figure 23: Results for the coefficients $B$ that quantifies the $\cos(2\phi^{*})$ modulation.
@@ V.1 Experimental systematic studies
The entire analysis was performed with several alternative sets of cuts and PID criteria. One study used a smaller range in track momenta and angles. Another ignored the heavy gas Cherenkov detector in the SHMS, which resulted in significant kaon contamination of the pion sample above momenta of 3 GeV. A third study used a lower aerogel threshold, with correspondingly larger kaon contamination subtraction. The luminosity and HMS acceptance were verified to be accurate to within 3% by comparison of elastic electron-proton measurements to a fit to global data Bosted and Christy (2008). The optical properties of the spectrometers were verified using the kinematic constraints of both $ep$ elastic scattering and exclusive pion production.
@@ V.3 Radiative corrections uncertainty
We have listed our results Hall C pion-SIDIS results both with and without radiative corrections, to allow for future improvements.
V.4 Interpretation systematics
Due to a lack of experimental data in our kinematic region, we have not corrected our results for the contributions from diffractive exclusive vector meson production. This contribution was studied by the COMPASS collaboration Agarwala et al. (2020), which found substantial corrections to $\left<cos(\phi^{*})\right$> and $\left<\cos(2\phi^{*})\right$> for $x<0.05$, but very small corrections for $x>0.1$, where all of the present data lie.
The data in this paper were taken at a single beam energy, and thus cannot be used to separate the transverse and longitudinal structure functions.
@@ VI Summary
In summary, we have measured the $\pi^{\pm}$ multiplicities from SIDIS on H and D targets over a range of $z$, $P_{t}$, and $\phi^{*}$. The multiplicities were fit for each bin in $(x,~Q^{2},~z,~P_{t})$ with three parameters: $\phi^{*}$ independent $M_{0}$ and azimuthal modulations $\langle\cos(\phi^{*})\rangle$ and $\langle\cos(2\phi^{*})\rangle$. The $P_{t}$-dependence of the $M_{0}$ results was found to be remarkably consistent for the four flavor cases studied: $ep\rightarrow e\pi^{+}X$, $ep\rightarrow e\pi^{-}X$, $ed\rightarrow e\pi^{+}X$, $ed\rightarrow e\pi^{-}X$ over the range $0<P_{t}<0.4$ GeV, as were the multiplicities evaluated near $\phi^{*}=180^{\circ}$ over the extended range $0<P_{t}<0.7$ GeV. Gaussian widths of the $P_{t}$-dependence exhibit a quadratic increase with $z$. The $\cos(\phi^{*})$ modulations were found to be consistent with zero for $\pi^{+}$, in agreement with previous world data, while the $\pi^{-}$ moments are in many cases significantly greater than zero. The $\cos(2\phi^{*})$ modulations were found to be consistent with zero. The higher statistical precision of this dataset compared to previously published data should allow improved determinations of quark transverse momentum distributions and higher twist contributions.
@@ VII Acknowledgments
We are grateful to M. Cerutti for providing calculations of the 2022 MAPS model at each of the kinematic settings of this experiment

### v2

@@ IV.4 Pion multiplicities near $\phi^{*}=180^{\circ}$
Figure 14: The fit parameter $M_{0}$ weighted by $z^{2}$ as a function of $P_{t}$, averaged over $150<\phi^{*}<210$ degrees, for the three kinematic settings of this experiment (top to bottom) in four bins in $z$ (left to right). Within each panel, the results from top to bottom are for $\pi^{+}$ from a proton target (black), $\pi^{+}$ from a deuteron target (red), $\pi^{-}$ from a deuteron target (green) and $\pi^{-}$ from a proton target (blue). Relative systematic errors (not shown) are estimated to be 1.3% (2.2%) point-to-point for setting III (settings I and $I$) and 1.8% for the scale uncertainty. The curves are the predictions of the MAP collaboration based on a fit to previous world data MAPS; MAP24; MAP22_4, normalized with a parameter $k$ to give the best overall agreement with these results. The dashed curves illustrate two-parameter Gaussian fits to each dataset of the form $M=ae^{(-bP_{t}^{2})}$. No fully differential results from previously published experiments are available for comparison.
Due to the experimental setup limitations, it was not possible to obtain full azimuthal coverage at large $P_{t}$. Nonetheless, a considerable amount of time was spent accumulating data near $\phi^{*}=$180, which primarily measures $M_{0}(1-A+B)$. Given this caveat, it is of interest to plot the multiplicities at $\left<\phi^{*}\right$$\sim$ 180, over the full $P_{t}$ experimental range, as shown in Fig. 14. The MAP calculations for $M_{0}$ (MAP did not include non-zero values of A and B in their fits) describe the $P_{t}$-independence quite well up to $P_{t}=0.5$ GeV, which is the largest value provided by the collaboration. Although not perfect in magnitude, the Gaussian-shaped fits shown in the figure describe the $P_{t}$-independence remarkably well to values of $P_{t}$ as high as 0.7 GeV.
@@
Figure 16: Ratios to $(p,\pi^{+})$ of experimental multiplicities averaged over $150<\phi^{*}<210$ degrees for $(d,\pi^{+})$ (red), $(p,\pi^{-})$ (green), and $(d,\pi^{-})$ (blue). Systematic errors (not shown) are discussed in the text. The solid curves are the ratios of MAP calculations, while the dashed curves are the ratios of Gaussian fits from Fig. 14. The curves follow the same color scheme as the data points.
To further quantify the observed flavor independence of the multiplicities, the ratio of $\left<\phi^{*}\right$$\sim$ 180 multiplicities for $(d,\pi^{+})$, $(p,\pi^{-})$, and $(d,\pi^{-})$ to the multiplicities for $(p,\pi^{+})$ are plotted in Fig. 16, as a function of $P_{t}$ for the three kinematic settings and four $z$ bins. The ratios are consistent with no statistically significant dependence on $P_{t}$. The ratios are compared to those from the MAP calculations MAPS; MAP24; MAP22_4, which also show only very slight $P_{t}$-dependence.
@@ IV.5 Azimuthal dependence of multiplicities
The $\phi^{*}$ dependence of the measured multiplicities is quantified by the two coefficients, $A$ and $B$, associated with the $\cos(\phi^{*})$ and $\cos{(2\phi^{*})}$ modulations of the multiplicities. The $\cos(\phi^{*})$ coefficient, $A$, obtained from the fit of the multiplicity results in each $(z,P_{t})$ bin to functional form– Eq. 10, is shown in Fig. 17 as a function of $P_{t}$. These results show an overall trend that $A$ for $\pi^{-}$ production on both protons and deuterons is significantly >> 0 at high $z$ for all three kinematic settings. On the other hand, the $A$ coefficient is consistent with zero within experimental uncertainties or has small negative values for $\pi^{+}$ production on both protons and deuterons. The $\pi^{+}$ results are consistent with the previous HERMES measurements  HERMES:2012kpt but have the opposite sign for the $\pi^{-}$, except for the HERMES point HERMES:2012kpt at highest $x$. We note that all of these data generally lie at higher $x$ than those from HERMES HERMES:2012kpt or COMPASS COMPASS_PT_ref.
Figure 17: Results from this experiment (solid circles) for the coefficients AA that scale the $\cos(\phi^{*})$ distributions for $(p,\pi^{+})$ (black), $(d,\pi^{+})$ (red), $(p,\pi^{-})$ (green), and $(d,\pi^{-})$ (blue). Relative systematic errors (not shown) are estimated to be 0.01 (0.02) point-to-point for setting III (settings I and II). The scale uncertainty is negligibly small. The open circles are from HERMES at $Q^{2}=6.6$ GeV2, x=x=0.36 and $\left<z\right$>=0.45Barone:2015ksa, and curves are from a fit to HERMES and COMPASS data Barone:2015ksa; HERMES:2012kpt. The curves follow the same color scheme as the data points.
Similarly the $\cos{(2\phi^{*})}$ coefficient, $B$, obtained from the fit of the multiplicity results in each $(z,P_{t})$ bin to Eq. 10 is shown in Fig. 18 as a function of $P_{t}$. Other than a couple of $z$ bins, these results show small values of $B$, that are either consistent with zero or have small positive values. These results are consistent with the previous HERMES measurements HERMES:2012kpt except at $\left<z\right$>=0.55.
Figure 18: Results from this experiment (solid circles) for the coefficients $B$ that scale the $\cos(2\phi^{*})$ distributions for $(p,\pi^{+})$ (black), $(d,\pi^{+})$ (red), $(p,\pi^{-})$ (green), and $(d,\pi^{-})$ (blue). Relative systematic errors (not shown) are estimated to be 0.01 (0.02) point-to-point for setting III (settings I and II). The scale uncertainty is negligibly small. The open circles are from HERMES at $Q^{2}=6.6$ GeV2, xx=0.36 and $\left<z\right$>=0.45Barone:2015ksa, and curves are from a fit to HERMES and COMPASS data Barone:2015ksa; HERMES:2012kpt. The curves follow the same color scheme as the data points.
@@ V.1 Experimental systematic studies
The entire analysis was performed with several alternative sets of cuts and PID criteria. One study used a smaller range in track momenta and angles. Another ignored the heavy gas Cherenkov detector in the SHMS, which resulted in significant kaon contamination of the pion sample above momenta of 3 GeV. A third study used a lower aerogel threshold, with correspondingly larger kaon contamination subtraction. The luminosity and HMS acceptance were verified to be accurate to within 3% by comparison of elastic electron-proton measurements to a fit to global data bosted_fit. The optical properties of the spectrometers were verified using the kinematic constraints of both $ep$ elastic scattering and exclusive pion production.
@@ V.2 Experimental systematic errors
In ratios of $\pi^{+}/\pi^{-}$, most systematic errors cancel out except for the uncertainty in particle identification. In ratios of proton to deuteron target for a fixed flavor, most systematic errors cancel except for the relative target thickness.
@@ V.3 Radiative corrections uncertainty
We have listed our results webpage both with and without radiative corrections, so that future global analyses can improve on the cross section models by incorporating data in kinematic regions not constrained by the present experiment, as has been done previously for DIS data.
V.4 Relevance to TMD factorization
The authors of Ref. Boglione:2022gpv have presented three kinematic regions for pion electroproduction: the target fragmentation region; a soft central region; and a current fragmentation region where TMD factorization should be applicable. They postulate that the probability of the current fragmentation region dominating by more than $2\sigma$ requires that the hadron rapidity $y_{h}<-1$, corresponding the the regions to the left of the arrows in Fig. 12. Most of the present data lie in this region.
Some previous experiments HERMES:2012kpt; COMPASSrho have subtracted contributions from pions originating from diffractive $\rho$ electroproduction, a process that is clearly not easily accounted for in the TMD factorization framework. Due to a lack of experimental data in our kinematic region, we have not corrected our results for the contributions from diffractive vector meson production. Based on an extrapolation of the COMPASS collaboration fit COMPASSrho, there could substantial corrections to the $\left<cos(\phi^{*})\right$> and $\left<\cos(2\phi^{*})\right$> moments, but small corrections for $x>0.1$, where all of the present data lie.
The data presented in this paper were taken at a single beam energy, and thus cannot be used to separate the transverse and longitudinal structure functions.
@@ VI Summary
In summary, we have measured the $\pi^{\pm}$ multiplicities from SIDIS on H and D targets over a three-dimensional grid in $z$, $P_{t}$, and $\phi^{*}$ for several values of $(x,Q^{2}$). The multiplicities were fitted for each bin in $(x,~Q^{2},~z,~P_{t})$ with three parameters: $\phi^{*}$ -independent $M_{0}$, and azimuthal modulations $\langle\cos(\phi^{*})\rangle$ and $\langle\cos(2\phi^{*})\rangle$. The $P_{t}$-dependence of the $M_{0}$ results was found to be remarkably consistent for the four flavor cases studied: $ep\rightarrow e\pi^{+}X$, $ep\rightarrow e\pi^{-}X$, $ed\rightarrow e\pi^{+}X$, $ed\rightarrow e\pi^{-}X$ over the range $0<P_{t}<0.4$ GeV, as were the multiplicities evaluated near $\phi^{*}=180^{\circ}$ over the extended range $0<P_{t}<0.7$ GeV. Gaussian widths of the $P_{t}$-dependence exhibit a quadratic increase with $z$. The $\cos(\phi^{*})$ modulations were found to be consistent with zero for $\pi^{+}$, in agreement with previous world data, while the $\pi^{-}$ moments were in many cases significantly greater than zero. The $\cos(2\phi^{*})$ modulations were found to be consistent with zero. The higher statistical precision of this dataset compared to previously published data should allow improved determinations of quark transverse momentum distributions and higher twist contributions.
The spin-averaged SIDIS results of this paper are but one part of a larger program at JLab, which includes the use of nuclear targets, polarized beams and targets, and a range of beam energies to separate longitudinal and transverse structure functions.
@@ VII Acknowledgments
We are grateful to M. Cerutti for providing calculations of the 2022 MAP model at each of the kinematic settings of this experiment